OpenStudy (anonymous):

Roger Waters can scrape up just enough money to rent a canoe for 1 hour and 30 minutes. If he paddles out on a lake at 4 kph and returns at 2 kph, how far will he have paddled the canoe? Show work and table.

OpenStudy (dumbcow):

im guessing we are assuming he is constantly paddling. d = r*t however far out he went, he has to paddle the same distance to get back to shore 4t1 = 2t2 --> t2 = 2t1 this means it takes him twice as long to get back t1 + t2 = 1.5 by substitution t1 + 2t1 = 1.5 3t1 = 1.5 t1 = 0.5 t2 = 1 d = 4*0.5 = 2*1 = 2 he paddled 2 km out and 2 km back for total of 4 km